A car is travelling at a constant speed of 950m/s. Passes a trooper hidden behind a Billboard one second of the car passes the Billboard, the trooper sets out from the Billboard to catch it, acelerating at a constant speed of 100m/s².how long does it take her to take the car?
1
Expert's answer
2020-02-20T09:41:30-0500

Let the time of motion of the trooper is t, then distance traveled by her


"d=\frac{at^2}{2}."

The distance traveled by a car


"d=v(t+1)."

Hence


"\frac{at^2}{2}=v(t+1),""\frac{100t^2}{2}=950(t+1),""t^2-19t-19=0,""t=20\:\rm s."

Finally, the distance traveled by trooper


"d=\frac{at^2}{2}=\frac{100\times 20^2}{2}=20
\:000\:\rm m."